# Truth, Trust, and Trouble: Medical AI on the Edge

### v1

@@ Abstract
Large Language Models (LLMs) hold significant promise for transforming digital health by enabling automated medical question answering. However, ensuring these models meet critical industry standards for factual accuracy, usefulness, and safety remains a challenge, especially for open-source solutions. We present a rigorous benchmarking framework via a dataset of over 1,000 health questions. We assess model performance across honesty, helpfulness, and harmlessness. Our results highlight trade-offs between factual reliability and safety among evaluated models—Mistral-7B, BioMistral-7B-DARE, and AlpaCare-13B. AlpaCare-13B achieves the highest accuracy (91.7%) and harmlessness (0.92), while domain-specific tuning in BioMistral-7B-DARE boosts safety (0.90) despite smaller scale. Few-shot prompting improves accuracy from 78% to 85%, and all models show reduced helpfulness on complex queries, highlighting challenges in clinical QA.
@@
Mohammad Anas Azeez1, Rafiq Ali2, Ebad Shabbir2, Zohaib Hasan Siddiqui1, Gautam Siddharth Kashyap3, Jiechao Gao4thanks: Corresponding Author: jiechao@stanford.edu, Usman Naseem3 1Jamia Hamdard, New Delhi, India 2DSEU-Okhla, New Delhi, India 3Macquarie University, Sydney, Australia 4Center for SDGC, Stanford University, California, USA
@@ 1 Introduction
Large Language Models (LLMs) are rapidly transforming digital health applications, from symptom checking Gupta et al. (2025) to medical Q&A Li et al. (2023). However, aligning these models to key industry-aligned principles—honesty, helpfulness, and harmlessness—remains a critical challenge. While proprietary models like GPT-4 Chang (2023) and Claude 3.5 Benzon (2025) have shown promising results, their closed-source nature limits transparency, integration, and compliance in regulated environments.
@@ 3 Methodology
Form passages $\{P_{1},\dots,P_{M}\}$ by clustering $s$ with overlapping $E(s)$QA Generation:
@@ 3.1 Data Collection and QA Generation
We then generate True/False QA pairs through two mechanisms. First, a rule-based template engine =$\mathcal{H}=\{h_{1},\dots,h_{T}\}$ creates factual assertions (e.g., “The {ANATOMICAL_ENTITY} is part of the {REGION_NAME}”) using entity–region tuples $(e,r)$. Each candidate $q_{\mathrm{templ}}$ is labeled based on its presence in $P_{j}$ as according to Equation (1).
@@
Only validated pairs are retained. To enhance safety and robustness, we define a curated set of edge-case patterns =$\mathcal{E}=\{e_{1},\dots,e_{E}\}$ representing known unsafe practices. Each QA pair is scanned for such risks using Equation (3).
@@
To ensure the reliability and safety of our QA dataset, all True/False questions were manually reviewed by a team of three licensed medical annotators (i.e. the authors), each holding advanced degrees in clinical medicine and anatomy. Annotators independently labeled a subset of examples for correctness, safety, and factual consistency. Disagreements were resolved through majority voting. To quantify inter-annotator reliability, we computed Cohen’s Kappa on overlapping subsets, yielding an average agreement of $\kappa=0.81$, indicating substantial consensus. We began with a pool of approximately 1,500 candidate QA pairs, derived from both rule-based (750) and LLM-generated (750) pipelines. After quality assurance filtering and annotation validation, 1,077 examples were retained for downstream evaluation. These post-processed examples form the benchmark used in our real-world applicability study. To assess robustness, we conducted a double-blind A/B testing protocol on this finalized set of 1,077 medical questions. Annotators, blinded to model identity and prompt source, evaluated system-generated responses to mitigate confirmation and source bias. This protocol enabled unbiased performance comparisons between QA generation strategies. To address potential biases in the dataset itself, we ensured balanced coverage across anatomical regions, question types (e.g., compositional, causal, negation), and document sources. Additionally, we applied pattern-based safety filters and edge-case screening (see Section 3.1) to exclude QA pairs exhibiting potentially harmful, misleading, or ungrounded content. Note: The complete dataset, including all validated QA pairs and annotation metadata, will be publicly released upon acceptance.
@@ 4.2 Hyperparameters
Our fine-tuning approach for domain-specific models such as AlpaCare-13B, BioMistral-7B-DARE, and Mistral-7B employs low learning rates ($1\times 10^{-5}$ to $5\times 10^{-5}$), moderate batch sizes (16–32), and weight decay to prevent overfitting, with early stopping after 3–5 epochs based on validation loss. During inference, parameters including temperature (0.7), top-$p$ sampling (0.9), and maximum token length (128 tokens) are tuned to optimize response relevance, informativeness, and diversity. Safety filtering is enforced via strict regex and keyword pattern matching with high sensitivity thresholds, triggering exclusion or manual review of any flagged content to minimize unsafe outputs. Evaluation metrics for factual consistency, helpfulness, and harmlessness apply semantic similarity thresholds between 0.80 and 0.85 on domain-adapted embeddings, ensuring reliable and meaningful QA generation. Experiments utilize multi-GPU clusters (NVIDIA A100) to support scalable fine-tuning and prompt inference pipelines that align with industry throughput and latency requirements. Code will be released post-acceptance.
@@ 5.5 Edge Case Generalization
5.6 Few-shot Prompting vs. Zero-shot
Incorporating in-context examples through few-shot prompting has become a promising approach to enhance language model reliability. We compared zero-shot prompting, which relies solely on the query, against few-shot prompting that includes three illustrative TRUE/FALSE examples with explanations before each query. Evaluations on factual accuracy, honesty, and helpfulness metrics (Table 6) demonstrate that few-shot prompting consistently improves model performance. Notably, accuracy increased from 78% to 85%, honesty improved by 7 percentage points, and helpfulness saw a 5 point gain. These gains suggest that contextual priming reduces hallucinations and bolsters factual consistency, aligning with industry priorities for deploying dependable AI systems. Organizations aiming for trustworthy clinical decision support should consider integrating few-shot techniques to enhance transparency and reduce error rates without additional fine-tuning. Note: We did not include harmlessness in this evaluation, as the prompts were factual classification queries with minimal risk of eliciting harmful content, focusing instead on correctness and informativeness.
@@ 6 Conclusion
This study evaluated three clinical LLMs—Mistral-7B, BioMistral-7B-DARE, and AlpaCare-13B—on factual accuracy, safety, and reasoning. AlpaCare-13B achieved the best performance with an accuracy of 91.7% and a harmlessness score of 0.92, showcasing its effectiveness in clinical QA. BioMistral-7B-DARE, despite its smaller scale, attained a high safety score of 0.90, highlighting the benefits of domain-specific tuning. Few-shot prompting boosted accuracy from 78% to 85%.

### v2

@@ Abstract
Large Language Models (LLMs) hold significant promise for transforming digital health by enabling automated medical question answering. However, ensuring these models meet critical industry standards for factual accuracy, usefulness, and safety remains a challenge, especially for open-source solutions. We present a rigorous benchmarking framework via a dataset of over 1,000 health questions. We assess model performance across honesty, helpfulness, and harmlessness. Our results highlight trade-offs between factual reliability and safety among evaluated models—Mistral-7B, BioMistral-7B-DARE, and AlpaCare-13B. AlpaCare-13B achieves the highest accuracy (91.7%) and harmlessness (0.92), while domain-specific tuning in BioMistral-7B-DARE boosts safety (0.90) despite smaller scale. Few-shot prompting improves accuracy from 78% to 85%, and all models show reduced helpfulness on complex queries, highlighting challenges in clinical QA. Our code is available at: https://github.com/AnasAzeez/TTT
@@ 1 Introduction
Large Language Models (LLMs) are rapidly transforming digital health applications, from symptom checking Gupta et al. (2025) to medical Q&A Li et al. (2023). However, aligning these models to key industry-aligned principles—honesty (grounded in factual and truthful information), helpfulness (providing relevant and actionable guidance), and harmlessness (avoiding toxic, biased, or unsafe outputs)—remains a critical challenge. While proprietary models like GPT-4 Chang (2023) and Claude 3.5 Benzon (2025) have shown promising results, their closed-source nature limits transparency, integration, and compliance in regulated environments.
@@ 3 Methodology
Form passages $\{P_{1},\dots,P_{M}\}$ by clustering $s$ with overlapping E(s)E(s)QA Generation:
@@ 3.1 Data Collection and QA Generation
We then generate True/False QA pairs through two mechanisms. First, a rule-based template engine $\mathcal{H}=\{h_{1},\dots,h_{T}\}$ creates factual assertions (e.g., “The {ANATOMICAL_ENTITY} is part of the {REGION_NAME}”) using entity–region tuples $(e,r)$. Each candidate $q_{\mathrm{templ}}$ is labeled based on its presence in $P_{j}$ as according to Equation (1).
@@
Only validated pairs are retained. To enhance safety and robustness, we define a curated set of edge-case patterns $\mathcal{E}=\{e_{1},\dots,e_{E}\}$ representing known unsafe practices. Each QA pair is scanned for such risks using Equation (3).
@@
To ensure the reliability and safety of our QA dataset, all True/False questions were manually reviewed by a team of three licensed medical annotators (i.e. the authors), each holding advanced degrees in clinical medicine and anatomy. Annotators independently labeled a subset of examples for correctness, safety, and factual consistency. Disagreements were resolved through majority voting. To quantify inter-annotator reliability, we computed Cohen’s Kappa on overlapping subsets, yielding an average agreement of $\kappa=0.81$, indicating substantial consensus. We began with a pool of approximately 1,500 candidate QA pairs, derived from both rule-based (750) and LLM-generated (750) pipelines. After quality assurance filtering and annotation validation, 1,077 examples were retained for downstream evaluation. These post-processed examples form the benchmark used in our real-world applicability study. To assess robustness, we conducted a double-blind A/B testing protocol on this finalized set of 1,077 medical questions. Annotators, blinded to model identity and prompt source, evaluated system-generated responses to mitigate confirmation and source bias. This protocol enabled unbiased performance comparisons between QA generation strategies. To address potential biases in the dataset itself, we ensured balanced coverage across anatomical regions, question types (e.g., compositional, causal, negation), and document sources. Additionally, we applied pattern-based safety filters and edge-case screening (see Section 3.1) to exclude QA pairs exhibiting potentially harmful, misleading, or ungrounded content.
@@ 4.2 Hyperparameters
Our fine-tuning approach for domain-specific models such as AlpaCare-13B, BioMistral-7B-DARE, and Mistral-7B employs low learning rates ($1\times 10^{-5}$ to $5\times 10^{-5}$), moderate batch sizes (16–32), and weight decay to prevent overfitting, with early stopping after 3–5 epochs based on validation loss. During inference, parameters including temperature (0.7), top-$p$ sampling (0.9), and maximum token length (128 tokens) are tuned to optimize response relevance, informativeness, and diversity. Safety filtering is enforced via strict regex and keyword pattern matching with high sensitivity thresholds, triggering exclusion or manual review of any flagged content to minimize unsafe outputs. Evaluation metrics for factual consistency, helpfulness, and harmlessness apply semantic similarity thresholds between 0.80 and 0.85 on domain-adapted embeddings, ensuring reliable and meaningful QA generation. Experiments utilize multi-GPU clusters (NVIDIA A100) to support scalable fine-tuning and prompt inference pipelines that align with industry throughput and latency requirements.
@@ 5.5 Edge Case Generalization
5.6 Few-shot Prompting vs. Zero-shot
Incorporating in-context examples through few-shot prompting has become a promising approach to enhance language model reliability. We compared zero-shot prompting, which relies solely on the query, against few-shot prompting that includes three illustrative TRUE/FALSE examples with explanations before each query. Evaluations on factual accuracy, honesty, and helpfulness metrics (Table 6) demonstrate that few-shot prompting consistently improves model performance. Notably, accuracy increased from 78% to 85%, honesty improved by 7%, and helpfulness saw a 5% gain. These gains suggest that contextual priming reduces hallucinations and bolsters factual consistency, aligning with industry priorities for deploying dependable AI systems. Organizations aiming for trustworthy clinical decision support should consider integrating few-shot techniques to enhance transparency and reduce error rates without additional fine-tuning. Note: We did not include harmlessness in this evaluation, as the prompts were factual classification queries with minimal risk of eliciting harmful content, focusing instead on correctness and informativeness.
@@
(a) Semantic Distribution via t‑SNE
(b) Lexical Heatmap of Question Vocabulary
(c) Temporal Dynamics of Correctness and Difficulty
(d) TRUE Answer Word Cloud
(e) FALSE Answer Word Cloud
(f) Answer Type Evolution Over Time
Figure 2: Semantic and temporal analysis of question-answer behavior. (a) t-SNE shows semantic clustering with difficulty overlay. (b) Heatmap illustrates lexical distribution across question indices. (c) Rolling correctness vs. difficulty trends. (d–e) Word clouds highlight frequent terms in TRUE and FALSE answers. (f) Evolution of answer types over time.
Qualitative Analysis.
Figure 2 offers an in-depth qualitative assessment of the dataset’s semantic and behavioral characteristics. Subfigure (a) uses t-SNE to project the semantic space of question embeddings, revealing dense clustering patterns correlated with difficulty levels—indicating that harder questions tend to occupy semantically distinct regions. The lexical heatmap in (b) highlights word frequency across the question index, showing that specific anatomical terms dominate and vary with question position. Subfigure (c) illustrates temporal dynamics by plotting rolling correctness rates alongside difficulty, uncovering periodic dips in performance that align with more complex or ambiguous question segments. Word clouds in (d) and (e) differentiate lexical emphasis in TRUE and FALSE answers, with TRUE responses focusing on terms like “radius”, “clavicle”, and “scapula”, while FALSE answers include distractors such as “joint”, “breast”, and “border”. Finally, (f) tracks the evolution of answer types over time, showing non-random fluctuations between TRUE and FALSE labels—suggesting shifts in dataset reasoning demands or structural design. Collectively, these visualizations provide insights into the semantic structure, linguistic patterns, and temporal answer behaviors that shape model performance.
@@ 6 Conclusion
This study evaluated three clinical LLMs—Mistral-7B, BioMistral-7B-DARE, and AlpaCare-13B—on factual accuracy, safety, and reasoning. AlpaCare-13B achieved the best performance with an accuracy of 91.7% and a harmlessness score of 0.92, showcasing its effectiveness in clinical QA. BioMistral-7B-DARE, despite its smaller scale, attained a high safety score of 0.90, highlighting the benefits of domain-specific tuning. Few-shot prompting boosted accuracy from 78% to 85%. However, all models exhibited limitations on complex reasoning tasks. These results emphasize persistent challenges in clinical LLMs and the necessity of balancing accuracy, safety, and reasoning for real-world deployment.